OpenStudy (dancergirl45):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 A (1, 12) B (−1, −12) C (12, −1) D (12, 1)

pooja195 (pooja195):

Ok Equation 1 = x − 2y = 14 Equation 2= x + 3y = 9 x−2y=14 for x: x−2y+2y=14+2y (Add 2y to both sides) x=2y+14

pooja195 (pooja195):

Now we can plug that into E2 \[\huge~\rm~\bf 2y+14 + 3y = 9\] solve for y

OpenStudy (dancergirl45):

5Y+14=9

pooja195 (pooja195):

good! now subtract 14 from both sides 5y=9-14 what is 9-14=?

OpenStudy (dancergirl45):

-5

pooja195 (pooja195):

good now divide both sides by 5 5y/5=-5/5 -5/5= ?

OpenStudy (dancergirl45):

-1

pooja195 (pooja195):

good so y=-1

OpenStudy (dancergirl45):

ok and the other is -12

pooja195 (pooja195):

x=2(-1)+14 x=-2+14 x=12

OpenStudy (dancergirl45):

so c again

pooja195 (pooja195):

The answer wa obvious there was only on option that has -1 as the y value so yes it is C :) good job

OpenStudy (dancergirl45):

Thank you so much again! i got a 100%

pooja195 (pooja195):

np =)